# devoted to memory of Alexei Alexandrovich Starobinsky

### v15

@@ Abstract
We consider the local thermodynamics of the de Sitter state. The local temperature, which is the same for all points of the de Sitter space, is $T=H/\pi$, where $H$ is the Hubble parameter. It is twice larger than the Gibbons-Hawking temperature of the cosmological horizon, $T_{\rm GH}=H/2\pi$. The local temperature is not related to the cosmological horizon. It determines the rate of the activation processes, which are possible in the de Sitter environment. The typical example is the process of the ionization of the atom in the de Sitter environment, which rate is determined by temperature $T=H/\pi$. The local temperature determines the local entropy of the de Sitter vacuum state, and this allows to calculate the total entropy inside the cosmological horizon. The result reproduces the Gibbons-Hawking area law, which is related to the cosmological horizon, $S_{\rm hor}=4\pi KA$, where $K=1/16\pi G$. We extend the consideration of the local thermodynamics of the de Sitter state using the $f({\cal R})$ gravity. In this thermodynamics, the Ricci scalar curvature ${\cal R}$ and the effective gravitational coupling, $K$, are thermodynamically conjugate variables. The holographic connection between the bulk entropy of the Hubble volume, and the surface entropy of the cosmological horizon remains the same, $S_{\rm hor}=4\pi KA$, but where the gravitational coupling is $K=df/d{\cal R}$. The local temperature of the de Sitter vacuum suggests that it serves as the heat bath for the matter, and thus the de Sitter state is locally unstable towards the creation of matter and its further heating. The decay of the de Sitter vacuum due to such processes determines the quantum breaking time of the space-times with positive cosmological constant. We also consider the thermodynamics of de Sitter in the frame of the multi-metric gravity ensemble, where the heat exchange between different ”sub-Universes” in the ensemble leads to the common de Sitter expansion with common temperature $T=H/\pi$. Application of the local thermodynamics to the entropy of the Schwarzschild black hole is also considered.
@@ I Introduction
We consider the local thermodynamics of the de Sitter state of the expansion of the Universe. The term “local” means that we consider the de Sitter vacuum as the thermal state, which is characterized by the local temperature. This consideration is based on observation, that matter immersed in the de Sitter vacuum feels this vacuum as the heat bath with the local temperature $T=H/\pi$, where $H$ is the Hubble parameter. This temperature is twice larger than the Gibbons-Hawking one, and it has no relation to the cosmological horizon. The existence of the local temperature, suggests the existence of the other local thermodynamic quantities, which participate in the local thermodynamics of the de Sitter state. In addition to the the local entropy density $s$ and local vacuum energy density $\epsilon$, there are also the local thermodynamic variables related to the gravitational degrees of freedom.
@@
Although the local temperature is twice larger than the Gibbons-Hawking temperature assigned to the horizon, $T_{\rm GH}=H/2\pi$, there is the certain connection between the local thermodynamics and the thermodynamics of the event horizon. It appears that the total entropy of the volume $V_{H}$ bounded by the cosmological horizon coincides with the Gibbons-Hawking entropy, $S_{\rm bulk}=sV_{H}=A/4G=S_{\rm hor}$. This demonstrates that the local thermodynamics in the $(3+1)$ de Sitter is consistent with the global thermodynamics assigned to the cosmological horizon, although the origin of such bulk-surface correspondence is not very clear.
@@
We tested the local thermodynamics and the bulk-surface correspondence in the extension of general relativity – the $f({\cal R})$ gravity. The $f({\cal R})$ gravity in terms of the Ricci scalar ${\cal R}$ is one of the simplest geometrical models, which describes the dark energy and de Sitter expansion of the Universe.Starobinsky1980 ; Starobinsky2007 ; Felice2010 ; Clifton2012 ; Nojiri2010wj ; Nojiri2017ncd ; Odintsov2023b It was used to construct an inflationary model of the early Universe – the Starobinsky inflation, which is controlled by the ${\cal R}^{2}$ contribution to the effective action. This class of models, $f({\cal R})\propto{\cal R}-{\cal R}^{2}/M^{2}$, was also reproduced in the so-called $q$-theory,KlinkhamerVolovik2008d ; KlinkhamerVolovik2008c where $q$ is the 4-form field introduced by HawkingHawking1984 for the phenomenological description of the physics of the deep (ultraviolet) vacuum (here the sign convention for ${\cal R}$ is opposite to that in Ref. Starobinsky2007 ). The Starobinsky model is in good agreement with the observations. However, despite the observational success, the theory of Starobinsky inflation is still phenomenological. Due to a rather small mass scale $M$ compared with the Planck scale it is difficult to embed the model into a UV complete theory.Chaichian2023 ; Brinkmann2023 ; Lust2023
@@
The $f({\cal R})$ theory demonstrates that the effective gravitational coupling $K$ (it is the inverse Newton constant, $K=1/16\pi G$) and the scalar curvature ${\cal R}$ are connected by equation $K=df/d{\cal R}$. This suggests that $K$ and ${\cal R}$ are the thermodynamically conjugate variables.Volovik2023a ; Volovik2023b This pair of the gravitational variables is similar to the pair of the electrodynamic variables, electric field ${\bf E}$ and electric induction ${\bf D}$, which participate in the thermodynamics of dielectrics, and to the pair of magnetic thermodynamic variables, ${\bf B}$ and ${\bf H}$. Using the local thermodynamics with $T=H/\pi$, we obtained the general result for the total entropy of the Hubble volume, $S_{\rm bulk}=sV_{H}=4\pi KA=S_{\rm hor}$, where $K=df/d{\cal R}$ is the effective gravitational coupling and ${\cal R}=-12H^{2}$.
@@
In Sec. II we consider the elements of the local thermodynamics of the de Sitter state. It is based on the observation that the de Sitter vacuum serves as the thermal bath for the external matter – the matter immersed into the de Sitter environment. This is different from the traditional consideration of the vacua of the quantum fields in the de Sitter spacetime, which uses the Euclidean action method. Example of the influence of the de Sitter vacuum to the external matter is provided by an atom in the de Sitter environment (Sec. II.1). As distinct from the atom in the flat space, the atom in the de Sitter vacuum has a certain probability of ionization. The rate of ionization is similar to the rate of ionization in the presence of the thermal bath with temperature $T=H/\pi$.Volovik2009a ; Volovik2009 ; Volovik2023c ; Volovik2024c ; Maxfield2022 The same temperature determines the other activation processes, which are energetically forbidden in the Minkowski spacetime, but are allowed in the de Sitter background, see also Refs. Maldacena2015 ; Reece2023 (Sec. II.2). In Sec. II.3 it is shown that the local temperature also determines the Hawking temperature of radiation from the cosmological horizon without using the Euclidean action.
@@
All that suggests that it is natural to consider the temperature $T=H/\pi$ as the local temperature of the de Sitter vacuum. The local temperature leads to the local entropy of the de Sitter thermal bath (Sec. II.8), which integrated over the Hubble volume reproduces the Gibbons-Hawking entropy of the cosmological horizon (Sec.II.9). The effect of thermal fluctuations is discussed in Sec. III.1.
@@
Sec. IV is devoted to the multi-metric gravity, which can be considered as ensemble of the sub-Universes, each described by its own metric $g_{\mu\nu(n)}$ (or tetrads $e^{a}_{\mu(n)}$), by its own gravitational coupling $K_{n}$ and cosmological constant $\Lambda_{n}$. The heat exchange between the sub-Universes leads to their thermalization – the formation of the Universe in which the sub-Universes have the common Hubble parameter and thus the common temperature.
In Sec. V the local entropy of the de Sitter state allows to consider the thermodynamics of the Schwarzschild black hole, assuming that the black hole can be formed by relaxation of the gravastar (Sec. V.3) with the de Sitter core. Here the interior of the gravastar is represented by the contracting de Sitter state (Sec. V.2), which has negative Hubble parameter, and thus the negative temperature $T=H/\pi<0$ and negative entropy. Since the cosmological and black hole horizons cancels each other, the gravastar has zero entropy. In the process of relaxation of the gravastar to the black hole the de Sitter core with its negative entropy shrinks, which results in the Hawking-Bekenstein entropy of the black hole horizon (Sec. V.4). The entropy of the white hole is considered in Sec. V.5, and the heat exchange between black holes in the multi-metric gravity is in Sec. V.6.
@@
Sec. VI is devoted to thermodynamics of de Sitter in the $f({\cal R})$ gravity. It contains the pair of the thermodynamically conjugate variables, $K$ and ${\cal R}$ (Sec. VI.1). These variables together with the local temperature and local entropy provides the Gibbs-Duhem relation (Sec. VI.2). The general result for the total entropy of the Hubble volume, $S_{\rm bulk}=sV_{H}=4\pi KA=S_{\rm hor}$ is obtained in Sec. VI.3. This is considered on example of quadratic gravity in Sec. VI.4.
The Sec. VII is devoted to the de Sitter decay due to the thermalization of matter by the de Sitter heat bath, and finally the conclusion is in Sec. VIII.
II Thermodynamics of the de Sitter state
@@ II.1 Atom in de Sitter environment as thermometer
We consider the de Sitter thermodynamics using the Painlevé-Gullstrand (PG) form,Painleve ; Gullstrand where the metric in the de Sitter expansion is
$ds^{2}=-dt^{2}+(dr-v(r)dt)^{2}+r^{2}d\Omega^{2}\,.$ (1)
@@
Here the shift velocity is $v(r)=Hr$. This metric is stationary, i.e. does not depend on time, and it does not have the unphysical singularity at the cosmological horizon. That is why it is appropriate for consideration of the local thermodynamics both inside and outside the horizon. It also allows to describe two different states – the expanding de Sitter Universe with $H>0$ and contracting de Sitter Universe with $H<0$.
@@
Now let us consider an atom at the origin, $r=0$. The atom is the external object in the de Sitter spacetime, which is playing the role of the detector (or the role of the static observer) in this spacetime. The electron bounded to an atom may absorb the energy from the gravitational field of the de Sitter background and escape from the electric potential barrier. If the ionization potential is much smaller than the electron mass but is much larger than the Hubble parameter, $H\ll\epsilon_{0}\ll m$, one can use the nonrelativistic quantum mechanics to estimate the tunneling rate through the barrier. The corresponding radial trajectory $p_{r}(r)$ is obtained from the classical equation
$\frac{p_{r}^{2}}{2m}+p_{r}v(r)=-\epsilon_{0}\,,$ (2)
where $p_{r}(r)v(r)$ is the Doppler shift:
$p_{r}(r)=-mv(r)+\sqrt{m^{2}v^{2}(r)-2m\epsilon_{0}}\,.$ (3)
@@
The integral of $p_{r}(r)$ over the classically forbidden region, $0<r<r_{0}=\sqrt{2\epsilon_{0}/mH^{2}}$, gives the ionization rate
$w\sim\exp{\left(-2\,{\rm Im}\,S\right)}=\exp\left(-\frac{\pi\epsilon_{0}}{H}\right)\,.$ (4)
This is equivalent to the thermal thermal radiation with temperature $T=H/\pi$, which is twice the Gibbons-Hawking temperature of the cosmological horizon, $T=2T_{\rm GH}$.
@@
In the interpretation of Ref.Maxfield2022 , the bound state decays by quantum tunnelling from the point $r=0$ to the point $r=r_{0}$, at which the energy $-\epsilon_{0}$ matches the de Sitter gravitational potential $U(r)=-mH^{2}r^{2}/2$. The radial trajectory $p_{r}(r)$ is obtained from the classical equation
$\frac{{\bf p}^{2}}{2m}-\frac{1}{2}mH^{2}r^{2}=-\epsilon_{0}\,.$ (5)
@@
This gives Eq.(4) for the WKB tunneling rate.

### v16

# Thermodynamics and decay of de Sitter vacuum

@@ Abstract
We discuss the consequences of the unique symmetry of the de Sitter spacetime. This symmetry leads to the specific thermodynamic properties of the de Sitter vacuum, which produces the thermal bath for matter. The de Sitter spacetime is invariant under the modified translations, ${\bf r}\rightarrow{\bf r}-e^{Ht}{\bf a}$, where $H$ is the Hubble parameter. For $H\rightarrow 0$, this symmetry corresponds to the conventional invariance of Minkowski spacetime under translations ${\bf r}\rightarrow{\bf r}-{\bf a}$. Due to this symmetry, all the comoving observers at any point of the de Sitter space perceive the de Sitter environment as the thermal bath with temperature $T=H/\pi$, which is twice larger than the Gibbons-Hawking temperature of the cosmological horizon. This temperature does not violate the de Sitter symmetry, and thus does not require the preferred reference frame, as distinct from thermal state of matter, which violates the de Sitter symmetry. This leads to the heat exchange between gravity and matter, and to instability of de Sitter state towards the creation of matter, its further heating, and finally to the decay of the de Sitter state. The temperature $T=H/\pi$ determines different processes in the de Sitter environment, which are not possible in the Minkowski vacuum, such as the process of ionization of an atom in the de Sitter environment. This temperature also determines the local entropy of the de Sitter vacuum state, and this allows us to calculate the total entropy inside the cosmological horizon. The result reproduces the Gibbons-Hawking area law, which is related to the cosmological horizon, $S_{\rm hor}=4\pi KA$, where $K=1/(16\pi G)$. This supports the holographic properties of the cosmological event horizon. We extend the consideration of the local thermodynamics of the de Sitter state using the $f({\cal R})$ gravity. In this thermodynamics, the Ricci scalar curvature ${\cal R}$ and the effective gravitational coupling $K$ are thermodynamically conjugate variables. The holographic connection between the bulk entropy of the Hubble volume and the surface entropy of the cosmological horizon remains the same, but with the gravitational coupling $K=df/d{\cal R}$. Such connection takes place only in the $3+1$ spacetime, where there is the special symmetry due to which the variables $K$ and ${\cal R}$ have the same dimensionality. We also consider the lessons from the de Sitter symmetry for the thermodynamics of black and white holes.
@@ I Introduction
We consider the local thermodynamics of the de Sitter stage of the expansion of the Universe. The term “local” means that we consider the de Sitter vacuum as the thermal state, which is characterized by the local temperature. This consideration is based on observation, that matter immersed in the de Sitter vacuum feels this vacuum as the heat bath with the local temperature $T=H/\pi$, where $H$ is the Hubble parameter. This temperature has no relation to the cosmological horizon, and to the Hawking radiation from the cosmological horizon. However, it is exactly twice the Gibbons-Hawking temperature, $T_{\rm GH}=H/2\pi$. The reason for such relation is the symmetry of the de Sitter space-time with respect to the combined translations. In the Minkowski vacuum, this symmetry becomes the conventional invariance of under translations.
The existence of the local temperature suggests the existence of the other local thermodynamic variables, which participate in the local thermodynamics of the de Sitter state. In addition to the the local entropy density $s$ and local vacuum energy density $\epsilon$, there are also the local thermodynamic variables related to the gravitational degrees of freedom.
@@
Although the local temperature is twice larger than the Gibbons-Hawking temperature assigned to the cosmological horizon, there is the certain connection between the local entropy $s$ and the global entropy usually assigned to the event horizon. It appears that the total entropy of the de Sitter state in the volume $V_{H}$ bounded by the cosmological horizon coincides with the Gibbons-Hawking entropy, which is proportional to the area $A$ of the horizon, $sV_{H}=A/4G$. Such holographic bulk-surface correspondence takes place only in the $3+1$ spacetime. Although the peculiarity of the $3+1$ spacetime may be related to the special symmetry, which connects the local thermodynamic variables, the origin of the holography is still not very clear. That is why we extended the consideration of the local thermodynamics to the $f({\cal R})$ gravity, and checked the bulk-surface correspondence in this modified version of general relativity.
@@
The $f({\cal R})$ gravity in terms of the Ricci scalar ${\cal R}$ is one of the simplest geometrical models, which describes the dark energy and de Sitter expansion of the Universe.Starobinsky1980 ; Starobinsky2007 ; Felice2010 ; Clifton2012 ; Nojiri2010wj ; Nojiri2017ncd ; Odintsov2023b It was used to construct an inflationary model of the early Universe – the Starobinsky inflation, which is controlled by the ${\cal R}^{2}$ contribution to the effective action. This class of models, $f({\cal R})\propto{\cal R}-{\cal R}^{2}/M^{2}$, was also reproduced in the so-called $q$-theory,KlinkhamerVolovik2008d ; KlinkhamerVolovik2008c where $q$ is the 4-form field introduced by HawkingHawking1984 for the phenomenological description of the physics of the deep (ultraviolet) vacuum (here the sign convention for ${\cal R}$ is opposite to that in Ref. Starobinsky2007 ). The Starobinsky model is in good agreement with the observations. However, despite the observational success, the theory of Starobinsky inflation is still phenomenological. Due to a rather small mass scale $M$ compared with the Planck scale it is difficult to embed the model into a UV complete theory.Chaichian2023 ; Brinkmann2023 ; Lust2023 ; Elizalde:2017mrn ; Odintsov:2017hbk ; Myrzakulov:2014hca ; Bamba:2014jia ; Sebastiani:2013eqa ; Ketov2024 But we used this model only for the generalization of the de Sitter thermodynamics and for consideration of the validity of the holographic principle.
@@
The $f({\cal R})$ theory demonstrates that the effective gravitational coupling $K$ (it is the inverse Newton constant, $K=1/16\pi G$) and the scalar curvature ${\cal R}$ are connected by equation $K=df/d{\cal R}$. This suggests that $K$ and ${\cal R}$ are the thermodynamically conjugate variables.Volovik2023a ; Volovik2023b This pair of the non-extensive gravitational variables is similar to the pair of the electrodynamic variables, electric field ${\bf E}$ and electric induction ${\bf D}$, which participate in the thermodynamics of dielectrics. They are also similar to the pair of magnetic thermodynamic variables, magnetic induction ${\bf B}$ and magnetic field ${\bf H}$.
In the de Sitter spacetime, the local temperature does no depend on the theories of general relativity and thus has the same value $T=H/\pi$. Using the local thermodynamics with this temperature, we obtained the general result for the total entropy of the Hubble volume, $S_{\rm bulk}=sV_{H}=4\pi KA=A/4G=S_{\rm hor}$, where $K=df/d{\cal R}$ is the effective gravitational coupling and ${\cal R}=-12H^{2}$. This supports the holographic bulk-surface correspondence in the $3+1$ spacetime.
Note that the $3+1$ spacetime has special symmetry, which is absent in the other dimensions. The thermodynamically conjugate variables have the same dimensionality: all of them have dimensionality of the square of the inverse length: $[K]=[{\cal R}]=[1/l^{2}]$, $[{\bf B}]=[{\bf H}]=[1/l^{2}]$ and $[{\bf D}]=[{\bf E}]=[1/l^{2}]$. It looks that this symmetry is important for the validity of the holographic correspondence.
Since the de Sitter state has thermal behaviour, it serves as the thermal bath for matter. The matter violates the de Sitter symmetry, as a result the energy exchange between vacuum and matter leads to the decay of the de Sitter state.
The plan of the paper is the following.
@@
In Sec. II we show that the de Sitter vacuum serves as the thermal bath for matter immersed in the de Sitter environment. Our approach differs from the traditional consideration of the vacua of the quantum fields in the de Sitter spacetime, which uses the Euclidean action method. An example of the influence of the de Sitter vacuum on immersed matter is an atom in the de Sitter environment (Sec. II.1). As distinct from the atom in the flat space, the atom in the de Sitter vacuum has a certain probability of ionization. The rate of ionization is similar to the rate of ionization of an atom in the flat space-time in the presence of the thermal bath with temperature $T=H/\pi$.Volovik2009a ; Volovik2009 ; Volovik2023c ; Volovik2024c ; Maxfield2022 In Sec. II.2 it is shown that the same temperature determines the other activation processes, which are energetically forbidden in the Minkowski spacetime, but are allowed in the de Sitter background. Examples are the splitting of the heavy particle with mass $m$ to two particles with masses $m_{1}+m_{2}>m$,Maldacena2015 ; Reece2023 and radiation of electron-positron pairs by the electron at rest. In Sec. II.4 it is shown that the local temperature $T$ also determines the Gibbons-Hawking temperature of radiation from the cosmological horizon $T_{\rm GH}=T/2$ without using the Euclidean action.
@@
Sec. III is devoted to the local thermodynamics of the de Sitter state, which is determined by the local temperature $T=H/\pi$. The local temperature leads to the local entropy of the de Sitter thermal bath (Sec. III.2), which integrated over the Hubble volume reproduces the Gibbons-Hawking entropy of the cosmological horizon (Sec. III.4).
@@
Sec. IV describes the attempt to obtain the de Sitter thermal states from the general principles of thermodynamics of the many-body systems. We consider the multi-metric gravity, which can be viewed as an ensemble of the sub-Universes, each being described by its own metric $g_{\mu\nu(n)}$ (or tetrads $e^{a}_{\mu(n)}$), by its own gravitational coupling $K_{n}$ and cosmological constant $\Lambda_{n}$. The heat exchange between the sub-Universes leads to their thermalization – the formation of the Universe in which the sub-Universes have the common Hubble parameter and thus they have the common temperature.
@@
Sec. V is devoted to thermodynamics of de Sitter in the $f({\cal R})$ gravity. The $f({\cal R})$ gravity contains the pair of the thermodynamically conjugate variables, $K$ and ${\cal R}$ (Sec. V.1). These variables together with the local temperature and local entropy provide the generalization of Gibbs-Duhem relation for the de Sitter state (Sec. V.2). The confirmation of the holographic result for the total entropy of the Hubble volume in the $f({\cal R})$ gravity is obtained in Sec. V.3: $S_{\rm bulk}=sV_{H}=4\pi KA=S_{\rm hor}$, The quadratic gravity and its symmetry are discussed in Sec. V.4.
The Sec. VI is devoted to the de Sitter decay due to the thermalization of matter by the de Sitter heat bath, and by the thermal fluctuations of the de Sitter state. These two mechanisms lead to different power laws of the decay, which may correspond to two different epochs.
Sections VII and VIII demonstrate how the local entropy of the de Sitter state allows us to consider the thermodynamics of the Schwarzschild black hole. The starting point of our consideration in Sec. VII.3 is that the black hole can be obtained from the relaxation of the gravastar object – the black hole, which has the de Sitter core. It is important that the de Sitter interior of the gravastar is represented by the contracting de Sitter state. The contracting de Sitter in Sec. VII.2 has negative Hubble parameter $H<0$, and thus the negative temperature $T=H/\pi<0$ and the negative entropy. The contracting and expanding de Sitter states can be considered as two phases obtained from the symmetric state of the Minkowski vacuum by the spontaneous breaking of the time reversal symmetry. These phases transform to each other under the time reversal, $t\rightarrow-t$. In this sense, the Hubble parameter $H$ can be considered as the order parameter of the symmetry breaking phase transition.
Since in the considered gravastar object, the cosmological horizon and the black hole horizon cancel each other, the gravastar has zero entropy. In the process of relaxation of the gravastar to the black hole, the de Sitter core with its negative entropy shrinks. This results in the Hawking-Bekenstein entropy of the black hole horizon (Sec. VII.4). The negative entropy of the white hole is considered in Sec. VII.5. The heat exchange between black holes in the multi-metric gravity is discussed in Sec. VII.8. Section VIII provides the alternative derivation of the thermodynamics of black and white holes using the macroscopic quantum tunneling.
Finally the conclusion is in Sec. IX.
II de Sitter state as heat bath for matter
@@ II.1 Atom in de Sitter environment as thermometer
We consider the de Sitter thermodynamics using the Painlevé-Gullstrand (PG) form,Painleve ; Gullstrand where the metric is
$ds^{2}=-dt^{2}+(d{\bf r}-{\bf v}({\bf r})dt)^{2}\,.$ (1)
Here ${\bf v}({\bf r})$ is the shift velocity, which in condensed matter plays the role of the velocity of the inviscid superfluid component of the liquid (the velocity of the ”superfluid quantum vacuum” Volovik2003 ). In the de Sitter expansion the velocity of the ”vacuum” is ${\bf v}({\bf r})=H{\bf r}$ and the metric is
$ds^{2}=-dt^{2}+(dr-Hrdt)^{2}+r^{2}d\Omega^{2}\,.$ (2)
@@
The PG metric is stationary, i.e. does not depend on time, and it does not have the unphysical singularity at the cosmological horizon. That is why it is appropriate for consideration of the local thermodynamics both inside and outside the horizon. It also allows us to consider two different phases of the vacuum with broken time reversal symmetry. These are the expanding de Sitter Universe with $H>0$ and the contracting de Sitter Universe with $H<0$. These two phases transform to each other under the time reversal, $t\rightarrow-t$. In this sense, the Hubble parameter $H$ can be considered as the order parameter of the symmetry breaking phase transition from the symmetric state of the Minkowski vacuum.
@@
Now let us consider an atom at the origin, $r=0$. The atom is the external object in the de Sitter spacetime, which violates the de Sitter symmetry. It plays the role of the detector (or the role of the static observer) in this spacetime. The electron bounded to an atom may absorb the energy from the gravitational field of the de Sitter background and escape from the electric potential barrier. If the ionization potential is much smaller than the electron mass but is much larger than the Hubble parameter, $H\ll\epsilon_{0}\ll m$, one can use the nonrelativistic quantum mechanics to estimate the tunneling rate through the barrier.
Let us consider an electron on the $n$-th level in the hydrogen atom. Under the de Sitter gravitational field this electron can escape from the atom with the conservation of energy, which in the classical limit is given by the classical equation:
$\frac{p_{r}^{2}}{2m}+p_{r}v(r)=-E_{n}\,\,,\,\,E_{n}=\frac{me^{4}}{2\hbar^{2}}\frac{1}{n^{2}}\,.$ (3)
Here $p_{r}$ is the momentum of electron in the radial direction, $v(r)=Hr$ and $p_{r}v(r)$ is the Doppler shift, which allows for electron to reach the negative energy $-E_{n}$ when it escapes from the atom. The corresponding radial trajectory $p_{r}(r)$ for the escape of an electron from the atom is:
$p_{r}(r)=-mv(r)+\sqrt{m^{2}v^{2}(r)-2mE_{n}}\,.$ (4)
@@
The integral of $p_{r}(r)$ over the classically forbidden region, $0<r<r_{n}=\sqrt{2E_{n}/mH^{2}}$, gives the ionization rate in the semiclassical WKB approximpation:
$w\sim\exp{\left(-2\,{\rm Im}\,S\right)}=$
[…]
This ionization rate is equivalent to the rate of ionization in the flat Minkowski spacetime in the presence of the heat bath with temperature $T=H/\pi$.
This heat bath temperature is twice the Gibbons-Hawking temperature $T_{\rm GH}$ usually attributed to the cosmological horizon, $T=2T_{\rm GH}$. Moreover, the electron trajectory is well inside the horizon, since $r_{n}\ll r_{H}=1/H$, and thus this process has no relation to the Hawking radiation. However, in Sec. II.4 we show that the relation between the temperature of ionization and Gibbons Hawking temperature, $T=2T_{\rm GH}$, is the property of the de Sitter symmetry.
@@
Since the process of ionization takes place well inside the cosmological horizon, one can use the conventional static metric with the gravitational potential $U(r)=-mH^{2}r^{2}/2$.Maxfield2022 Then the bound state decays by quantum tunnelling from the point $r=0$ to the point $r=r_{n}$, at which the electron level $-E_{n}$ matches the de Sitter gravitational potential $U(r_{n})=-mH^{2}r_{n}^{2}/2$. The radial trajectory $p_{r}(r)$ is obtained from the classical equation
$\frac{{\bf p}^{2}}{2m}-\frac{1}{2}mH^{2}r^{2}=-E_{n}\,,$ (6)
which gives for the radial trajectory:
$p_{r}(r)=\sqrt{m^{2}H^{2}r^{2}-2mE_{n}}\,.$ (7)
@@
The imaginary part of the action at this trajectory again gives Eq.(5) for the WKB tunneling rate.